Suppose x and y satisfy both of follwing equations below. Then y = ?
x-3y=7 and 2x+4y=5
Thanks.

Answer by adamchapman(301)

x-3y=7 ..1
2x+4y=5 ..2
Rearrange equation (..1):
x=3y+7 ..3
Substitute equation (..3) into equation (..2) for x:
2(3y+7)+4y=5
6y+14+4y=5
10y+14=5
10y=-9
y=-0.9 ..4
Substitute equation (..4) into equation (..3) for y:
x=-2.7+7
x=4.3
I hope this helps.
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would send me your feedback on my help to adam.chapman@student.manchester.ac.uk